# Variable Elliptic Structures on the Plane: Transport Dynamics, Rigidity, and Function Theory

### v4

@@
(“Shut up and calculate!” — David Mermin)
@@ Introduction
This expression appears universally once the structure varies. It measures the incompatibility between differentiation in the base variables and multiplication in the fibers. As a consequence, it simultaneously governs the inhomogeneity of generalized Cauchy–Riemann systems and the failure of the Leibniz rule for the natural first–order operators associated with the variable structure. These phenomena are not imposed by analytic assumptions; they are forced by the geometry of the moving algebra itself.
@@
selects exactly the conservative case of the transport law. In this regime, the Burgers equation becomes unforced, and several independent obstructions collapse simultaneously: the generalized Cauchy–Riemann operator becomes a derivation, the real Cauchy–Riemann system becomes homogeneous, and the generator $i(x,y)$ is transported compatibly with its own multiplication law. Rigidity is therefore not the source of the coefficient dynamics, but a compatibility condition that restores integrability within an already dynamical setting.
The explicit $\varepsilon$–family of rigid elliptic structures constructed in this work by standard techniques provides concrete examples of this conservative regime. These examples show that rigidity does not force triviality: genuinely nonconstant structures exist, although ellipticity naturally restricts the domain of definition. From the transport viewpoint, these restrictions reflect characteristic phenomena familiar from first–order hyperbolic systems.
@@ 1.6 Explicit Formula for the Inhomogeneous Coefficients
1.7 Product Rule and Its Defect
Let $f=u+vi$ and $g=p+qi$ be $C^{1}$ sections. Define the Leibniz defect
$\mathcal{N}(f,g):=\partial_{\bar{z}}(fg)-(\partial_{\bar{z}}f)g-f(\partial_{\bar{z}}g).$
A direct computation shows
$\mathcal{N}(f,g)=\frac{1}{2}\,vq\,(i_{x}+i\,i_{y}).$ (1.13)
Remark 1.4.
Formula (1.13) is purely algebraic and holds without further assumptions. It shows explicitly how the variability of the generator $i(x,y)$ obstructs the derivation property of $\partial_{\bar{z}}$.
@@ 2.5 Universal Transport Law
2.6 Characteristic Interpretation
Equation (2.6) may be written in transport form as
$(\partial_{x}+\lambda\,\partial_{y})\lambda=G_{0}+\lambda\,G_{1}.$
Along characteristic curves satisfying $dy/dx=\lambda$, the spectral parameter evolves according to the linear ordinary differential equation
$\frac{d\lambda}{dx}=G_{0}(x,y(x))+\lambda(x)\,G_{1}(x,y(x)).$
Thus, the evolution of the coefficients of the structure reduction is always governed by transport along characteristic curves. The obstruction $\mathcal{G}$ acts as a source and growth term.
@@ 2.7 Summary
Chapter 3 Rigidity as the Conservative Regime of Burgers Transport
@@ 3.4 Rigidity and Conservative Transport
3.5 Collapse of Obstructions
We now show that rigidity causes several independent obstructions to collapse simultaneously.
3.5.1 Leibniz defect collapse
Let $f=u+vi$ and $g=p+qi$ be $C^{1}$ sections. As shown in Chapter 1, the Leibniz defect of the generalized Cauchy–Riemann operator is given by
$\mathcal{N}(f,g)=\frac{1}{2}\,vq\,\mathcal{G}.$
Proposition 3.4 (Leibniz defect collapse).
The operator $\partial_{\bar{z}}$ satisfies the Leibniz rule on $C^{1}(\Omega;\allowbreak A)$ if and only if the structure is rigid.
Proof.
The defect vanishes for all $f,g$ if and only if $\mathcal{G}\equiv 0$. ∎
3.5.2 Inhomogeneity collapse
@@ 3.6 The Rigidity Theorem
    The generalized Cauchy–Riemann operator satisfies the Leibniz rule.
  4. 4.
    The real Cauchy–Riemann system is homogeneous.
@@
The equivalence $(1)\Leftrightarrow(2)$ was shown above. Equivalences with $(3)$ and $(4)$ follow from the collapse results. ∎
@@ 3.7 Interpretation
In chapter 5 we show that this conservative regime is precisely what allows integral representation formulas and a coherent function theory to re-emerge.
Chapter 4 Conservative Transport and Explicit Rigid Structures
@@ 4.2 The Burgers Substrate
4.3 Recalling the Characteristic Formulation
We utilize the standard method of characteristics to generate solutions. Equation (4.1) implies that $\lambda$ is constant along the integral curves of the vector field $(1,\lambda)$:
$\frac{dx}{ds}=1,\qquad\frac{dy}{ds}=\lambda,\qquad\frac{d\lambda}{ds}=0.$
This classical result immediately implies the properties relevant to our construction:
  • Propagation: Structural data is transported linearly along lines in the complexified $(x,y)$ space.
  • Implicit Solutions: The general solution is given implicitly by $\lambda=F(y-\lambda x)$ for an arbitrary function $F$.
  • Blow–up: Smooth solutions generally develop singularities in finite time (characteristic crossing), which corresponds to the degeneration of the elliptic structure.
In the following sections, we apply this standard machinery to the specific ansatz of linear dependence, recovering a canonical family of rigid structures.
@@ 4.4 Conservative Burgers Transport
4.5 Characteristic Formulation
Equation (4.1) admits the transport form
$(\partial_{x}+\lambda\,\partial_{y})\lambda=0.$ (4.3)
Thus $\lambda$ is constant along characteristic curves $(x(s),y(s))$ satisfying
$\frac{dx}{ds}=1,\qquad\frac{dy}{ds}=\lambda,\qquad\frac{d\lambda}{ds}=0.$
Equivalently, along integral curves of the vector field $(1,\lambda)$ in the (x,y)(x,y)–plane, the spectral parameter is conserved. This immediately implies:
  • local existence for arbitrary $C^{1}$ initial data;
  • breakdown through characteristic crossing;
  • strong constraints on global smooth solutions.
These are dynamical consequences of the transport structure and do not depend on ellipticity.
@@ 4.8 Reduction to Ordinary Differential Equations
Let
$a=a(x),\qquad b=b_{0}>0\ \text{constant}.$
Substituting $a=a(x)$ and $b=b_{0}$ (constant) into (4.4), the second equation is identically satisfied, while the first reduces to
$a_{x}+a\,a_{y}=0.$
Since $a=a(x)$, this forces $a_{y}=0$, so the system imposes no further restriction on $a(x)$ at this stage. To obtain a nontrivial rigid structure, it is more transparent to work directly in terms of $(\alpha,\beta)$.
The nontrivial case is obtained by allowing $a_{x}\neq 0$ while keeping $b$ constant, which forces linear dependence in $y$.
@@
Because the rigidity condition is most naturally expressed in the original coefficients $(\alpha,\beta)$, we now translate the ansatz back into these variables.
@@ 4.12 Interpretation
This family serves as the canonical explicit model throughout the remainder of the theory.
@@
Chapter 5 The Cauchy–Pompeiu Formula in the Rigid Regime
@@ 5.3 A multiplicative norm
Since the algebra is commutative, the ii–terms cancel, and using $i^{2}=-\beta i-\alpha$ gives $v^{2}i^{2}=v^{2}(\beta i+\alpha)$; the ii–part cancels, leaving $u^{2}+\beta uv+\alpha v^{2}\in\mathbb{R}$. Multiplicativity follows from $\widehat{ab}=\hat{a}\,\hat{b}$:
@@
$Q(z)=\begin{pmatrix}1&\beta(z)/2\\ \beta(z)/2&\alpha(z)\end{pmatrix},\qquad(u,v)Q(z)(u,v)^{T}=N_{z}(u+vi(z)).$
@@ 5.5 $\partial_{\bar{z}}$-holomorphicity of the kernel under rigidity
On $\Omega\setminus\{\zeta\}$ we have $Z\cdot Z^{-1}=1$. By the Leibniz rule for $\partial_{\bar{z}}$ (rigidity),
@@ 5.8 Cauchy–Pompeiu Representation Theorem
Next compute $\partial_{\bar{z}}g=\partial_{\bar{z}}(fZ^{-1})$. Under rigidity, $\partial_{\bar{z}}$ satisfies Leibniz, so
@@
Chapter 6 Covariant Holomorphicity and Gauge Structure in the Rigid Regime
@@ 6.1 Motivation from the Cauchy–Pompeiu Formula
Chapter 5 established that under ellipticity and rigidity, every $f\in C^{1}(\Omega;A)$ satisfies the Cauchy–Pompeiu representation
@@ 6.3 Basic Algebraic Properties
Under rigidity, $\partial_{\bar{z}}$ satisfies the Leibniz rule
@@
Thus even in the rigid regime, covariantly holomorphic functions are not closed under the usual pointwise product. In order to remedy this will introduce a scalar gauge transformation.
@@ 6.9.4 Step 3: compute the right–hand side and compare
Chapter 7 Power Series Representations in the Rigid Regime
7.1 Purpose and Overview
In the classical theory of complex analysis, holomorphic functions admit local power series representations obtained by differentiating the Cauchy integral formula. This procedure relies crucially on two structural facts:
  • the imaginary unit is constant,
  • the Cauchy kernel depends only on the difference $z-\zeta$.
Neither of these properties survives in the presence of a variable elliptic structure. As a result, differentiation under the integral sign is no longer a legitimate operation.
The purpose of this chapter is twofold:
  1. 1.
    to show that covariantly holomorphic functions admit intrinsic power series representations in frozen algebra powers;
  2. 2.
    to explain how ordinary holomorphic functions inherit power series expansions only via gauge trivialization, and not directly.
This resolves the apparent tension between the failure of classical arguments and the persistence of a local expansion theory.
Important scope distinction. Throughout this chapter, the term “power series representation” refers to expansions constructed from integral representations and valid on non-infinitesimal neighborhoods. Purely infinitesimal Taylor expansions at a point, which rely only on freezing the algebra at that point, do not detect the variable structure and may exist without invoking any gauge. The necessity of gauge trivialization arises only when transport or integral manipulations are involved.
7.2 Standing Assumptions
Throughout this chapter we assume:
  • $\alpha,\beta\in C^{1}(\Omega)$ with $\Delta=4\alpha-\beta^{2}>0$;
  • the structure is rigid, i.e.
    $i_{x}+i\,i_{y}=0;$
  • $A_{z}=\mathbb{R}[X]/(X^{2}+\beta(z)X+\alpha(z))$;
  • $Df=\bar{\partial}f+\tfrac{1}{2}f\,i_{y}$ is the covariant Cauchy–Riemann operator.
All statements are local.
7.3 Covariantly Holomorphic Functions and the Cauchy–Pompeiu Formula
Recall that in the rigid regime the Cauchy–Pompeiu representation takes the form
$f(\zeta)=\frac{1}{2\pi j(\zeta)}\int_{\partial\Omega}^{(\zeta)}\frac{f(z)}{Z(z,\zeta)}\,d\tilde{z}-\frac{1}{\pi j(\zeta)}\iint_{\Omega}^{(\zeta)}\frac{\bar{\partial}f+\tfrac{1}{2}f\,i_{y}}{Z(z,\zeta)}\,dx\,dy.$
Lemma 7.1.
If $f\in C^{1}(\Omega;A)$ satisfies $Df=0$, then
$f(\zeta)=\frac{1}{2\pi j(\zeta)}\int_{\partial\Omega}^{(\zeta)}\frac{f(z)}{Z(z,\zeta)}\,d\tilde{z}$.
Proof.
This follows immediately from the vanishing of the interior term $\bar{\partial}f+\tfrac{1}{2}f\,i_{y}$ in the Cauchy–Pompeiu formula. ∎
This boundary representation is the correct replacement for the classical Cauchy integral formula in the variable elliptic setting.
7.4 Frozen Algebra and Polynomial Powers
Fix $\zeta=(\xi,\eta)\in\Omega$ and define the frozen generator $i(\zeta)\in A_{\zeta}$. Introduce the frozen variable
$Z_{0}(z,\zeta):=(y-\eta)-i(\zeta)(x-\xi)\in A_{\zeta}.$
All powers $Z_{0}^{n}$ are computed in the fixed algebra $A_{\zeta}$. They are genuine polynomials in $(x-\xi,y-\eta)$ and form a linearly independent family over $\mathbb{R}$.
The variable kernel $Z(z,\zeta)$ differs from $Z_{0}(z,\zeta)$ only through the dependence of $i(z)$ on $z$, which is controlled by ellipticity and rigidity.
7.5 Expansion of the Cauchy Kernel
On sufficiently small circles $|z-\zeta|=r$, the inverse kernel admits the convergent expansion
$\frac{1}{Z(z,\zeta)}=\sum_{n=0}^{\infty}\frac{Z_{0}(z,\zeta)^{n}}{\rho(z,\zeta)^{n+1}},\qquad\rho(z,\zeta):=|Z_{0}(z,\zeta)|_{\zeta}^{2}.$
This expansion is purely algebraic and takes place entirely in the frozen fiber $A_{\zeta}$. Uniform ellipticity guarantees convergence.
7.6 Power Series for Covariantly Holomorphic Functions
Theorem 7.2 (Power series for covariantly holomorphic functions).
Let $f\in C^{1}(\Omega;A)$ satisfy $Df=0$. Then for every $\zeta\in\Omega$ there exists $r>0$ and unique coefficients $a_{n}(\zeta)\in A_{\zeta}$ such that
$f(z)=\sum_{n=0}^{\infty}a_{n}(\zeta)\,Z_{0}(z,\zeta)^{n},\qquad|z-\zeta|<r,$
with uniform convergence on compact subsets. These coefficients arise from the Cauchy–Pompeiu representation and are therefore intrinsic to the covariant structure.
Proof.
Fix $\zeta$ and apply the boundary representation on $\partial B_{r}(\zeta)$. Insert the kernel expansion and integrate term-by-term. Coefficientwise integration is legitimate since all terms lie in the fixed algebra $A_{\zeta}$.
Define
$a_{n}(\zeta):=\frac{1}{2\pi j(\zeta)}\int_{\partial B_{r}(\zeta)}^{(\zeta)}\frac{f(z)}{\rho(z,\zeta)^{n+1}}\,d\tilde{z}$.
Ellipticity provides absolute convergence. Uniqueness follows from the linear independence of $\{Z_{0}^{n}\}$ in $A_{\zeta}$. ∎
7.7 Why Differentiation Under the Integral Fails in the Variable Case
We now explain why the classical argument based on differentiating the Cauchy integral formula does not extend to variable elliptic structures.
7.7.1 Dependence of the Kernel on the Evaluation Point
In the classical case, the kernel $(z-\zeta)^{-1}$ depends on $\zeta$ only through translation. In the variable case, the kernel
$Z(z,\zeta)=(y-\eta)-i(z)(x-\xi)$
depends on $\zeta$ in two independent ways:
  • explicitly through $(\xi,\eta)$,
  • implicitly through the evaluation fiber $A_{\zeta}$.
Differentiating with respect to $\zeta$ therefore produces terms involving derivatives of the generator $i(z)$ transported into a different fiber. There is no canonical identification that allows these terms to be controlled.
7.7.2 Moving Fibers and Loss of Algebraic Compatibility
The integral
$\int^{(\zeta)}\frac{f(z)}{Z(z,\zeta)}\,d\tilde{z}$
takes values in the frozen fiber $A_{\zeta}$ by coefficientwise projection. Differentiating with respect to $\zeta$ changes the target algebra itself.
Thus differentiation and integration do not commute: the operation “differentiate in $\zeta$” is not fiberwise compatible with the coefficientwise integration used to define the integral.
7.7.3 Non-Closure of the Cauchy 1–Form
Even under rigidity, the Cauchy 1–form
$d\tilde{z}=dy-i(z)\,dx$
is not closed:
$d(d\tilde{z})=i_{y}\,dx\wedge dy.$
This produces the interior correction term in the Cauchy–Pompeiu formula and forces the introduction of the covariant operator $D$. Attempting to differentiate under the integral sign ignores this structural defect and leads to incorrect identities.
7.7.4 Conclusion
The failure of differentiation under the integral is not technical. It is structural and reflects:
  • the variability of the algebra,
  • the lack of a canonical global trivialization,
  • the non-closure of the Cauchy 1–form.
Power series expansions must therefore be constructed from the integral representation itself, not by differentiation.
Remark 7.3 (Infinitesimal versus transported expansions).
The frozen variable $Z_{0}(z,\zeta)$ captures the infinitesimal behavior of the algebra structure at the point $\zeta$. Any $C^{1}$ function admits a formal Taylor expansion in powers of $(x-\xi,y-\eta)$, and such expansions may be rewritten in terms of $Z_{0}$ without reference to the covariant structure.
However, these purely infinitesimal expansions do not interact correctly with integral representations or with transport away from the basepoint. The constructions in this chapter address power series derived from the Cauchy–Pompeiu formula, which necessarily encode non-infinitesimal behavior and therefore require the covariant framework.
7.8 Power Series for Holomorphic Functions via Gauge
Let $h\in C^{1}(\Omega;A)$ satisfy $\bar{\partial}h=0$. While $h$ admits a formal Taylor expansion at any point by freezing the algebra, such expansions do not respect the integral structure or transport properties of the variable elliptic system. To obtain power series representations compatible with the Cauchy–Pompeiu formula, a gauge trivialization is required.
Theorem 7.4 (Power series for holomorphic functions).
For every $\zeta\in\Omega$ there exist $r>0$ and coefficients $b_{n}(\zeta)\in A_{\zeta}$ such that
$h(z)=\sum_{n=0}^{\infty}b_{n}(\zeta)\,Z_{0}(z,\zeta)^{n},\qquad|z-\zeta|<r.$
Proof.
Expand $f$ by the previous theorem and absorb $\psi(\zeta)$ into the coefficients. Convergence is preserved since $\psi$ is smooth and nonvanishing. ∎
Remark 7.5.
The role of the gauge in this section is not to enable the existence of local Taylor jets, which are purely infinitesimal, but to reconcile ordinary holomorphicity with the covariant framework underlying integral representations. Without gauge trivialization, the notions of differentiation, integration, and algebra multiplication fail to commute in a structurally controlled way.
7.9 Final Remarks
Covariant holomorphicity is the correct analytic notion in variable elliptic structures. It admits an intrinsic power series theory without invoking forbidden operations.
Ordinary holomorphicity inherits power series expansions only through gauge trivialization, not directly.
This clarifies both the necessity and the strength of the covariant framework.
In summary, Taylor expansions probe only the frozen infinitesimal algebra at a point, whereas Cauchy–Pompeiu expansions probe the transport of the algebra itself. The latter, and not the former, necessitate the covariant and gauge-theoretic framework developed here.

### v5

@@
(“Shut up and calculate!” — N. David Mermin)
Preface to v5
This version corrects an error in the Leibniz defect formula of Chapter 1 (formerly §1.7): the operator $\partial_{\bar{z}}$ is a derivation for all variable elliptic structures, not only rigid ones. The Rigidity Theorem has been updated accordingly. Clarification between $p(x)$-analytics and $p$-analytics in the sense of Polozhii has been added. Chapter 3 (Algebra–Spectral Intertwining) is new. The Poincaré residual has been renamed to Beltrami residual.
March 2026
@@ Introduction
This expression appears universally once the structure varies. It measures the incompatibility between differentiation in the base variables and multiplication in the fibers. As a consequence, it governs the inhomogeneity of generalized Cauchy–Riemann systems for the natural first–order operators associated with the variable structure. This phenomenon is not imposed by analytic assumptions; it is forced by the geometry of the moving algebra itself.
@@
selects exactly the conservative case of the transport law. In this regime, the Burgers equation becomes unforced, and two independent obstructions collapse simultaneously: the real Cauchy–Riemann system becomes homogeneous, and the generator $i(x,y)$ is transported compatibly with its own multiplication law. Rigidity is therefore not the source of the coefficient dynamics, but a compatibility condition that restores integrability within an already dynamical setting.
The explicit $\varepsilon$–family of rigid elliptic structures constructed in this work by standard techniques provides concrete examples of this conservative regime. These examples show that rigidity does not force triviality: genuinely nonconstant structures exist, although ellipticity naturally restricts the domain of definition. From the transport viewpoint, these domain restrictions arise because the ellipticity condition $\Im\lambda>0$ imposes global constraints on the admissible coefficient data.
@@ 2.5 Universal Transport Law
2.6 Ellipticity of the Transport System
The transport equation (2.6) has the form
$(\partial_{x}+\lambda\,\partial_{y})\lambda=G_{0}+\lambda\,G_{1},$
where $\lambda=a+ib$ with $b>0$. It is natural to ask whether this equation admits a characteristic interpretation in $\mathbb{R}^{2}$.
Writing (2.6) as a real first–order system for $(a,b)$, the principal part in the yy–direction is governed by the coefficient matrix
$A=\begin{pmatrix}a&-b\\ b&a\end{pmatrix}$,
whose eigenvalues are $a\pm ib=\lambda$ and $\bar{\lambda}$. Since $b>0$, the eigenvalues are strictly complex. Consequently the real system is elliptic, not hyperbolic, and possesses no real characteristic curves.
This is not a defect of the formulation but a structural necessity: the ellipticity condition $4\alpha-\beta^{2}>0$ that defines the variable elliptic structure propagates directly into the ellipticity of the transport system governing its coefficients. The structure is elliptic at every level of the theory.
Remark 2.5 (Complex vector field).
The operator $\partial_{x}+\lambda\,\partial_{y}$ is a complex vector field on $\mathbb{R}^{2}$. Its ellipticity is equivalent to the statement that it has no real null curves, i.e. the equation $dy/dx=\lambda$ admits no real–valued solution. In the language of several complex variables, $\partial_{x}+\lambda\,\partial_{y}$ fails to be tangent to any real curve precisely because $\operatorname{Im}\lambda\neq 0$.
@@ 2.7 Summary
Chapter 3 The Algebra–Spectral Intertwining
The universal transport law of Chapter 2 governs the evolution of the spectral parameter $\lambda$. The Cauchy–Riemann operator $\partial_{\bar{z}}=\frac{1}{2}(\partial_{x}+i\,\partial_{y})$ of Chapter 1 governs the function theory at the algebra level. This chapter establishes the exact relationship between the two: the spectral map intertwines the algebra operator with the spectral transport operator, universally, for all variable elliptic structures.
3.1 The spectral map
Recall from Chapter 1 that the fiber algebra $A_{z}=\mathbb{R}[X]/(X^{2}+\beta X+\alpha)$ is generated over $\mathbb{R}$ by the moving generator $i=i(x,y)$ satisfying $i^{2}+\beta\,i+\alpha=0$. Under ellipticity ($4\alpha-\beta^{2}>0$), $A_{z}$ is isomorphic to $\mathbb{C}$ as a real algebra. The spectral parameter
$\lambda=\frac{-\beta+i\sqrt{4\alpha-\beta^{2}}}{2}$
is the image of the generator $i$ under this isomorphism.
Definition 3.1 (Spectral map).
The spectral map sends an AA–valued section $W=U+Vi$ ($U,V:\Omega\to\mathbb{R}$) to its complex–valued spectral image
$W_{\lambda}:=U+V\lambda\in\mathbb{C}.$
The spectral map is a pointwise real–algebra isomorphism. Two properties are immediate.
Proposition 3.2 (Injectivity).
Under ellipticity, $W_{\lambda}=0$ implies $W=0$.
Proof.
$U+V\lambda=0$ with $\operatorname{Im}\lambda>0$ gives $U=0$ and $V=0$. ∎
Proposition 3.3 (Ring homomorphism).
For any AA–valued sections $W=U+Vi$ and $Z=P+Qi$,
$(WZ)_{\lambda}=W_{\lambda}\cdot Z_{\lambda}.$
Proof.
The spectral map $i\mapsto\lambda$ is a real–algebra homomorphism on each fiber. ∎
3.2 The Leibniz rule for $\partial_{\bar{z}}$
Before establishing the intertwining, we show that the algebra–level operator is always a derivation.
Proposition 3.4 (Universal Leibniz rule).
The operator $\partial_{\bar{z}}=\frac{1}{2}(\partial_{x}+i\,\partial_{y})$ satisfies the Leibniz rule on $C^{1}(\Omega;A)$:
$\partial_{\bar{z}}(fg)=(\partial_{\bar{z}}f)\,g+f\,(\partial_{\bar{z}}g)$
for all AA–valued $C^{1}$ sections $f,g$. This holds for all variable elliptic structures.
Proof.
The structure relation $i^{2}+\beta\,i+\alpha=0$ is an identity of $C^{1}$ functions on $\Omega$. Differentiating with respect to $x$:
$2i\,i_{x}+\beta\,i_{x}+\alpha_{x}+\beta_{x}\,i=0,$
which is the identity $(2i+\beta)i_{x}=-\alpha_{x}-\beta_{x}\,i$. This is exactly the condition that $\partial_{x}(i^{2})=2i\cdot i_{x}$ holds in the fiber algebra. The same argument applies to $\partial_{y}$. Hence $\partial_{x}$ and $\partial_{y}$ are derivations on $C^{1}(\Omega;A)$.
Since the fiber algebra is commutative, any AA–linear combination of derivations is again a derivation. In particular, $\partial_{\bar{z}}=\frac{1}{2}(\partial_{x}+i\,\partial_{y})$ is a derivation. ∎
3.3 The universal spectral identity
Theorem 3.5 (Universal spectral identity).
For every AA–valued $C^{1}$ section $W=U+Vi$:
$2\bigl(\partial_{\bar{z}}W\bigr)_{\lambda}=(W_{\lambda})_{x}+\lambda\,(W_{\lambda})_{y}.$ (3.1)
This holds for all variable elliptic structures.
Proof.
From the algebra–level Cauchy–Riemann decomposition, writing $\mathcal{G}=G_{0}+G_{1}\,i$:
$2\,\partial_{\bar{z}}W=(U_{x}-\alpha V_{y}+VG_{0})+(V_{x}+U_{y}-\beta V_{y}+VG_{1})\,i.$
Its spectral image is
$2(\partial_{\bar{z}}W)_{\lambda}=(U_{x}-\alpha V_{y}+VG_{0})+\lambda(V_{x}+U_{y}-\beta V_{y}+VG_{1}).$ (3.2)
From $W_{\lambda}=U+V\lambda$, using $\lambda^{2}=-\beta\lambda-\alpha$ (the spectral image of the structure relation):
$(W_{\lambda})_{x}+\lambda(W_{\lambda})_{y}$ $=U_{x}+V_{x}\lambda+V\lambda_{x}+\lambda(U_{y}+V_{y}\lambda+V\lambda_{y})$
$=(U_{x}-\alpha V_{y})+(V_{x}+U_{y}-\beta V_{y})\lambda+V(\lambda_{x}+\lambda\lambda_{y}).$ (3.3)
The universal transport law (Chapter 2) gives $\lambda_{x}+\lambda\lambda_{y}=G_{0}+G_{1}\lambda$, so
$V(\lambda_{x}+\lambda\lambda_{y})=VG_{0}+VG_{1}\lambda.$
Substituting into (3.3) yields (3.2). ∎
Identity (3.1) says that the spectral map sends the algebra–level operator $\partial_{\bar{z}}$ to the first–order transport operator $\frac{1}{2}(\partial_{x}+\lambda\,\partial_{y})$:
$\begin{array}[]{ccc}C^{1}(\Omega;A)&\xrightarrow{\;\partial_{\bar{z}}\;}&C^{0}(\Omega;A)\\[3.0pt] \downarrow\scriptstyle{W\mapsto W_{\lambda}}&&\downarrow\scriptstyle{W\mapsto W_{\lambda}}\\[3.0pt] C^{1}(\Omega;\mathbb{C})&\xrightarrow{\;\frac{1}{2}(\partial_{x}+\lambda\partial_{y})\;}&C^{0}(\Omega;\mathbb{C})\end{array}$
Both horizontal arrows are derivations (the top by Proposition 3.4, the bottom because it is first–order), and the vertical arrows are ring homomorphisms (Proposition 3.3). The diagram commutes for all variable elliptic structures.
3.4 Consequences
Corollary 3.6 (Algebra holomorphicity implies spectral transport).
If $\partial_{\bar{z}}W=0$, then $(W_{\lambda})_{x}+\lambda(W_{\lambda})_{y}=0$.
Proof.
Immediate from (3.1): the left side vanishes. ∎
Corollary 3.7 (Kernel correspondence).
The spectral map restricts to a ring isomorphism between $\ker\partial_{\bar{z}}$ and $\ker\frac{1}{2}(\partial_{x}+\lambda\partial_{y})$.
Proof.
By the intertwining identity, $W$ is in the kernel of $\partial_{\bar{z}}$ if and only if $W_{\lambda}$ is in the kernel of $\frac{1}{2}(\partial_{x}+\lambda\partial_{y})$. Both maps preserve products (Proposition 3.3 and the Leibniz rules). ∎
3.5 Commutativity with higher derivatives
The spectral map also commutes with pure yy–derivatives.
Proposition 3.8.
For every AA–valued $C^{2}$ section $W=U+Vi$:
$(W_{yy})_{\lambda}=(W_{\lambda})_{yy}.$
The same holds for $\partial_{xx}$ and all higher pure partial derivatives.
Proof.
$W_{yy}=U_{yy}+V_{yy}\,i+2V_{y}\,i_{y}+V\,i_{yy}$. Under $i\mapsto\lambda$: $(W_{yy})_{\lambda}=U_{yy}+V_{yy}\lambda+2V_{y}\lambda_{y}+V\lambda_{yy}$. From $W_{\lambda}=U+V\lambda$: $(W_{\lambda})_{yy}=U_{yy}+V_{yy}\lambda+2V_{y}\lambda_{y}+V\lambda_{yy}$. The argument for $\partial_{xx}$ is identical. ∎
3.6 Interpretation
The results of this chapter are summarized as follows.
  1. 1.
    The algebra–level operator $\partial_{\bar{z}}$ is a derivation on $C^{1}(\Omega;A)$, universally (Proposition 3.4).
  2. 2.
    The spectral map $W\mapsto W_{\lambda}$ is a pointwise ring isomorphism (Proposition 3.3).
  3. 3.
    These two structures are compatible: the spectral map intertwines $\partial_{\bar{z}}$ with the first–order transport operator $\frac{1}{2}(\partial_{x}+\lambda\partial_{y})$ (Theorem 3.5).
  4. 4.
    The intertwining extends to second–order operators via commutativity with pure partial derivatives (Proposition 3.8).
The obstruction $\mathcal{G}=i_{x}+i\,i_{y}$ appears in the proof of the intertwining identity through the universal transport law, but it does not prevent any of these results from holding. Its role is to force the transport law and to make the Cauchy–Riemann system inhomogeneous—not to obstruct the algebraic or intertwining properties of $\partial_{\bar{z}}$.
Chapter 4 Rigidity as the Conservative Regime of Burgers Transport
@@ 4.4 Rigidity and Conservative Transport
4.5 Inhomogeneity collapse
@@ 4.6 The Rigidity Theorem
    The generalized real Cauchy–Riemann system is homogeneous.
@@
The equivalence $(1)\Leftrightarrow(2)$ was shown above. Equivalences with $(3)$ follow from the inhomogeneity collapse result. ∎
@@ 4.7 Interpretation
In chapter 6 we show that this conservative regime is precisely what allows integral representation formulas and a coherent function theory to re-emerge.
Chapter 5 Conservative Transport and Explicit Rigid Structures
@@ 5.10 Interpretation
This family serves as the first canonical explicit model throughout the remainder of the theory.
@@
Chapter 6 The Cauchy–Pompeiu Formula in the Rigid Regime
@@ 6.3 A multiplicative norm
Since the algebra is commutative, the ii–terms cancel, and using $i^{2}=-\beta i-\alpha$ gives $v^{2}i^{2}=v^{2}(\beta i+\alpha)$; the ii–part cancels, leaving $u^{2}-\beta uv+\alpha v^{2}\in\mathbb{R}$. Multiplicativity follows from $\widehat{ab}=\hat{a}\,\hat{b}$:
@@
$Q(z)=\begin{pmatrix}1&-\beta(z)/2\\ -\beta(z)/2&\alpha(z)\end{pmatrix},\qquad(u,v)Q(z)(u,v)^{T}=N_{z}(u+vi(z)).$
@@ 6.5 $\partial_{\bar{z}}$-holomorphicity of the kernel under rigidity
On $\Omega\setminus\{\zeta\}$ we have $Z\cdot Z^{-1}=1$. By the Leibniz rule for $\partial_{\bar{z}}$,
@@ 6.8 Cauchy–Pompeiu Representation Theorem
Next compute $\partial_{\bar{z}}g=\partial_{\bar{z}}(fZ^{-1})$. In general, $\partial_{\bar{z}}$ satisfies Leibniz, so
@@
Chapter 7 Covariant Holomorphicity and Gauge Structure in the Rigid Regime
@@ 7.1 Motivation from the Cauchy–Pompeiu Formula
Chapter 6 established that under ellipticity and rigidity, every $f\in C^{1}(\Omega;A)$ satisfies the Cauchy–Pompeiu representation
@@ 7.3 Basic Algebraic Properties
In general, $\partial_{\bar{z}}$ satisfies the Leibniz rule
@@
Thus covariantly holomorphic functions are not closed under the usual pointwise product. In order to remedy this will introduce a scalar gauge transformation.